OpenStudy (anonymous):

Carolyn owes $9,620 on a credit card with 16.4% interest rate. She owes $3,970 on another credit card that has a 24.6% interest rate. What is the total monthly payment needed to pay off both cards in 3 months, assuming she makes fixed payments and does not make anymore charges to her cards?

OpenStudy (ranga):

\[\Large M = \frac{ P(1 + r)^nr }{ (1 + r)^n - 1 }\] M is Monthly payment P is the amount borrowed r is interest rate per month (in decimals) n is the number of months of the loan

OpenStudy (ranga):

For the first credit card, P = 9620; r = 0.164/12; n = 3. Calculate M. Do similar calculations for the second credit card. Add the two monthly payments to get the total monthly payment.

OpenStudy (anonymous):

Okay, I did it. I got the answer! Thank you so much!

OpenStudy (ranga):

you are welcome.